OpenStudy (anonymous):

(sin x )sin x cos^2x = sin^3x

OpenStudy (anonymous):

prove identities

OpenStudy (anonymous):

are you sure it's an identity? \[\sin^{2}(x) *\cos(x)=\sin ^{3}(x)\]

OpenStudy (anonymous):

\[\sin x - \sin x \cos ^2x=\sin ^3x\]

OpenStudy (anonymous):

sorry for the confusion.

OpenStudy (anonymous):

okay, now we are talking :) \[=\sin(x) [1-\cos ^{2}(x)] \rightarrow \sin(x)*\sin ^{2}(x)= \sin ^{3}(x)\]

OpenStudy (anonymous):

ah! I see, because sin^2x+cos^2x=1 so sin^2x=1-cos^2x

OpenStudy (anonymous):

that's so right @Mathdude101 ;)